OpenStudy (anonymous):

PLEASE HELP! Perform the required operations on the following functions. Given: f(x) = 3 - x; g(x) = -2x Find g[f(x)].

OpenStudy (anonymous):

g[f(x)]=g(3-x) =-2(3-x) =2x-6